Chapter 2 Questions 1,2,3

PROBLEMS 1. Given the following demand schedule of a 4. commodity: P(S) 6 5 4 3 2 1 0 QD 0 10 20 30 40 50 60 Show that by substituting the prices given in the function, you obtain the corresponding quantities OD 60 10P table into the following demand equation or demanded given in the table: *2. Derive the demand schedule from the following demand function: OD 80- 10P 3. On the same graph, plot the demand schedule of Problem 1 and label it D and the demand curve of Problem 2 and label it D. Does D represent an increase in demand or an increase in the quantity demanded? Why?
Answer #1

1.

Demand Scheduled
Price QD = 60 - 10P Quantity
$ 6 60 - 10*6 = 60 - 60= 0 0
5 60 -10*5 = 60 - 50 = 10 10
4 60 - 10*4 = 60 - 40 = 20 20
3 60 - 10*3 = 60 - 30 = 30 30
2 60 - 10*2 = 60 - 20 = 40 40
1 60 - 10*1 = 60 - 10 = 50 50
0 60 - 10*0 = 60 60

2. Given, QD' = 80 - 10P

Demand Scheduled
Price Q = 80 - 10P Quantity
$ 8 80 - 10*8 = 0 0
7 80 - 10*7 = 10 10
6 80 - 10*6 = 20 20
5 80 - 10*5 = 30 30
4 80 - 10*4 = 40 40
3 80 - 10*3 = 50 50
2 80 - 10*2 = 60 60
1 80 -10*1 = 70 70
0 80 - 10*0 = 80 80

3. The graph I have plotted using excel kindly refer the attached picture

Demand Scheduled 100 80 60 C 40 20 0 4 8 -20 -40

Please contact if having any query thank you.